OpenStudy (anonymous):

In a bag, there are 2 blue tokens, 3 red tokens, 6 yellow tokens, and 4 green tokens. Stephen will perform random trials of drawing a token and replacing it into the bag. Which of the following is a good prediction for the number of times a red or green token will be drawn in 300 trials? A. 85 B. 119 C. 181 D. 138

OpenStudy (anonymous):

so there are 2+3+6+4 = 15 tokens in the bag, and out of those, 3 +4 =7 are red and green so there are 7/15 that you want to get. so for every pull there is just under half a chance that the one you pull out would be red or green, so half of 300 is 150, and 7/15 = ~.46 (just under half) so i would go with D, 138, as that is the only number close, but still below half. A is to small, B is smaller and less likely when compared to D, and it is not C as it has to be less then half of 300